OpenStudy (anonymous):

Ms.Suzuki invested $27,000 in two accounts, one yielding 4% interest and the other yielding 9%. If she recieved a total on $1,930 in interest at the end of the year, how much did she invest in each account?

hero (hero):

x + y = 27000 .04x + .09y = 1930

hero (hero):

Do you have any clue what to do from here?

OpenStudy (anonymous):

Well I would want to maybe use substitution to solve for one?

hero (hero):

yup, but...I recommend solving both equations for y first

OpenStudy (anonymous):

So, like, x+y=27000: Does y=27000-x?

hero (hero):

Yes, like that. Now do it for the other equation

OpenStudy (anonymous):

Does x=10,000?

OpenStudy (anonymous):

Not sure I came up with 10,000,17000

OpenStudy (anonymous):

Oh...wait, HA! That works! TY!

hero (hero):

you're welcome